# Assouad spectrum of Gatzouras–Lalley carpets

### v2

@@ Abstract
[…]
@@ 1.1 Overview
[…]
@@
Here, for a general set $E\subset\operatorname{{\mathbb{R}}}^{d}$ and $r>0$, $N_{r}(E)$ denotes the smallest number of balls $B(x,r)$ with $x\in E$ required to cover the set $E$. As $\theta$ converges to 00, $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ converges to the upper box dimension of $K$, and as $\theta$ converges to $1$, $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ converges to the quasi-Assouad dimension of Lü & Xi [zbl:1345.28019], as proven in [zbl:1410.28008]. In general, the quasi-Assouad dimension and Assouad dimension need not agree, though they do coincide for the sets under consideration in this document.
Beyond being a useful bi-Lipschitz invariant, the Assouad spectrum has also played an important role in applications: for instance, it plays a key role in the study of $L^{p}$-improving properties of spherical maximal operators with restricted dilation sets [zbl:07732556], and was used to prove sharp quasiconformal distortion estimates for polynomial spirals [zbl:07731255].
The Assouad spectrum has been explicitly computed for a wide variety of well-studied dynamical sets with distinct box and Assouad dimensions, such as some overlapping self-similar sets and self-affine sets [zbl:1407.28002, doi:10.54330/afm.142535], Bedford–McMullen carpets [zbl:1407.28002], certain Kleinian limit sets [zbl:1520.28005], parabolic Julia sets [arxiv:2203.04943], elliptical polynomial spirals [zbl:1510.28006], and random sets such as Mandelbrot percolation [zbl:1407.28002]. In all of these examples, the Assouad spectrum has a very particular form: it is piecewise convex, with pieces of the form $\theta\mapsto a+b\cdot\frac{\theta}{1-\theta}$ for some $a,b\in\operatorname{{\mathbb{R}}}$, and with points of non-differentiability corresponding to certain “geometrically meaningful” values. The Assouad spectrum has also been computed for certain infinitely generated self-conformal sets, where more complicated forms can appear based on the fine structure of the set of fixed points (which may be prescribed in a non-dynamical way) [arxiv:2207.11611].
@@
.47 001111001111 {subcaptionblock}.47 001111001111Refer to caption
@@
In this paper we focus on the problem of computing the Assouad spectrum for a particular class of self-affine sets, namely the Gatzouras–Lalley carpets first introduced in [zbl:0757.28011]. These generalise the well-studied Bedford–McMullen carpets, which were introduced in [Bedford] and [zbl:0539.28003]. The precise definition of the class of Gatzouras–Lalley carpets is deferred to Section 1.2, but a depiction of the rectangles representing the generating maps and the corresponding attractor for a specific example in this class is given in Fig. 3. Roughly speaking, the difference between Bedford–McMullen and Gatzouras–Lalley carpets is that all rectangles for a Bedford–McMullen carpet have the same widths and heights, while Gatzouras–Lalley carpets allow inhomogeneity across the widths of columns and also of the heights of rectangles within columns. However, we emphasize that the difference between Bedford–McMullen carpets and Gatzouras–Lalley carpets is more than just technical. For instance, it is known that Gatzouras–Lalley carpets need not have a unique measure of maximal dimension [zbl:1228.37025]. More famously, the simplest known example of a self-affine set with no ergodic measure of maximal dimension is defined using a 3-dimensional analogue of the Gatzouras–Lalley construction [zbl:1387.37026]. These features are not exhibited by Bedford–McMullen sponges in any dimension.
Our main contribution is to establish an explicit formula for the Assouad spectrum of a Gatzouras–Lalley carpet as the concave conjugate of the minimum of a finite family of elementary functions combined with a simple parameter change—see Section 1.3 for the precise statement. We observe novel and unexpected behaviour which has not been previously witnessed in any dynamically-invariant examples. More precisely, we obtain the following direct qualitative consequences of our explicit formula:
@@
    The Assouad spectrum can be a non-trivial differentiable function of $\theta$. In fact, we give a simple characterisation of differentiability, and observe that non-differentiability is generic (both topologically and measure theoretically) within the parameter space of Gatzouras–Lalley carpets—see Section 2.2.
@@
    The Assouad spectrum can have a non-trivial interval on which it is strictly concave. Again, this property holds generically—see Section 2.3
@@
    The Assouad spectrum is increasing and piecewise analytic, but with potentially arbitrarily many phase transitions and phase transitions of arbitrarily high odd order—see Remark 2.6.
These properties are direct consequences of the explicit formula, and their proofs can be found in Section 2. The proof of the explicit formula constitutes the majority of the work in this paper.
In contrast to the Bedford–McMullen case (for which the derivation of the Assouad spectrum is, generally speaking, straightforward), the Gatzouras–Lalley case has substantially more technical complications as a result of the inhomogeneity between columns and within each column. Our proof of the general formula uses recent techniques developed in the context of multifractal analysis, which we highlight here. First, in Section 3 we establish a general variational formula for the Assouad spectrum as a certain constrained maximisation problem involving information-theoretic quantities evaluated at Bernoulli measures. A key tool here is the method of types from large deviations theory, which was used in [zbl:07731132] to compute the $L^{q}$-spectrum of self-affine sponges and in [arxiv:2111.05625] to calculate the intermediate dimensions of Bedford–McMullen carpets. The explicit covering arguments build on and refine the fine covering strategies for self-affine carpets used in [zbl:1278.37032, zbl:1305.28021, arxiv:2309.11971]. Secondly, in LABEL:s:explicit we solve this variational formula to obtain our explicit formula. The main complexity here is that the variational formula is a non-smooth and non-convex optimisation problem. Our key technique here is the geometry of Lagrange duality, which was used in [arxiv:2312.08974] in order to elucidate the concave conjugate relationship apparent in the multifractal formalism.
@@ 1.2 Gatzouras–Lalley carpets and dimensions
Fix a finite index set $\mathcal{I}$ and for each $i\in\mathcal{I}$ fix constants $10<b_{i}<a_{i}<$ 1 and $c_{i},d_{i}\in\operatorname{{\mathbb{R}}}$. Consider the diagonal self-affine iterated function system (IFS) $\{T_{i}\}_{i\in\mathcal{I}}$ where $T_{i}\colon\operatorname{{\mathbb{R}}}^{2}\to\operatorname{{\mathbb{R}}}^{2}$ is given by
@@
As depicted in Fig. 3, the key features of a Gatzouras–Lalley IFS are that the rectangles $T_{i}([0,1]^{2})$ cannot overlap except possibly along edges, they lie in columns which themselves cannot overlap except possibly along edges, and the height of each rectangle is strictly less than its width.
@@
[…]
@@
We denote this quantity by $t_{\min}$ because of the usage in Eq. 2 below. We interpret $t_{\min}$ as the “average” column dimension, weighted appropriately using the column widths $a_{\underline{j}}$. Then the box dimension of $K$ was calculated in [zbl:0757.28011, Theorem 2.4] as
@@
$0\leq\min_{\underline{j}\in\eta(\mathcal{I})}s_{\underline{j}}\leq t_{\min}\leq\max_{\underline{j}\in\eta(\mathcal{I})}s_{\underline{j}}=t_{\max}\leq 1$
@@ 1.3 Main formula
Observe that $\tau$ is strictly increasing on $[t_{\min},t_{\max}]$ and $\tau(t_{\max})=0$ is the unique zero of $\tau$, with $\tau(t)<0$ for $t_{\min}\leq t<t_{\max}$. Moreover, $\tau$ is a minimum of concave functions, and is therefore concave. We denote its concave conjugate by
@@
In Section 2, we derive a more transparent formula for the Assouad spectrum, and deduce many interesting qualitative features of the spectrum from it. The proof of Section 1.3 is subsequently given in Section 3 and LABEL:s:explicit
@@ 2 Consequences of the main formula
Throughout this section, we recall the notation introduced in Section 1, in particular the notation used in the statement of Section 1.3.
@@ 2.1 An alternative formula for the Assouad spectrum
of the interval $[t_{\min},t_{\max}]$, with corresponding parts $I_{n}=[t_{n-1},t_{n}]$, such that for each $n\in\operatorname{{\mathbb{N}}}$, there is a $\underline{j}_{n}$ so that
@@
and moreover for all $11\leq n\leq m-$ 1, $\psi_{\underline{j}_{n}}\neq\psi_{\underline{j}_{n+1}}$. The latter property ensures that the partition $(I_{n})_{n=1}^{m}$ is uniquely determined. We refer to this partition as the spectrum partition associated with the IFS $\{T_{i}\}_{i\in\mathcal{I}}$. We associate with the spectrum partition the following additional information, all of which depends only on the underlying IFS.
@@
Using this notation, Section 1.3 yields the following explicit piecewise formula for the Assouad spectrum of $K$. For clarity in the below formula, note that $g$ is strictly increasing with $g(t_{\max})=0$. The cases of this formula are depicted in Fig. 4.
@@
We recall the definition of $\tau$ and the main result proven in Section 1.3. Note that the formulae for $\theta\leq\theta_{\min}$ and $\theta\geq\theta_{\max}$ follow directly by the definition of the concave conjugate applied at the endpoint. Similarly, for $\theta_{\min}<\theta<\theta_{\max}$, the formulae follow directly since
@@
The parts in case Item 2 correspond to the interiors of $I_{n}$, in which case $g^{*}$ and $g_{n}^{*}$ agree, and the parts in case Item 3 correspond to the points of non-differentiability of $g$, which can only occur on the endpoints between adjacent $I_{n}$. In the latter case, $g^{*}$ is an affine function of $\alpha$ with an explicit formula depending only on the value of $g_{n}$ at $t_{n}$.
@@
It is also clear directly from the definition that $\theta_{\min}=\theta_{\max}$ if and only if $g$ is a affine function on the interval $[t_{\min},t_{\max}]$, which occurs if and only if $m=1$ and $I_{1}$ is homogeneous.
We note that the formulae in Item 1 and Item 4 are in fact special cases of Item 3 after substituting the respective value of $t_{n}$. In these three cases, expanding the definition of $\phi$, the formula is of the form $a+b\frac{\theta}{1-\theta}$, which has occurred previously as discussed in the introduction. In contrast, case Item 2 is novel and very much not of this form. This case occurs only in the presence of an inhomogeneous column.
In case Item 2, an implicit formula for the concave conjugate can be obtained from LABEL:p:minimisers (also see the discussion in LABEL:r:optim). The spectrum can also be obtained parametrically as a function of $t$: given a part $I_{n}=[t_{n-1},t_{n}]$ with corresponding column function $\psi=\psi_{\underline{j}_{n}}=g_{n}$, the graph of the function $\theta\mapsto\operatorname{dim}^{\theta}_{\mathrm{A}}K$ on $(\theta_{n,\min},\theta_{n,\max})$ is the same as the image of the interval $(t_{n-1},t_{n})$ under the map
@@
A graphical depiction of the curve $g(t)$ is given in Fig. 4, and the decomposition provided by Section 2.1 is given in Fig. 5. The case $\theta_{\min}=\theta_{\max}$ is satisfied, for example, by every Bedford–McMullen carpet (or more generally by any Gatzouras–Lalley carpet with $\frac{\log b_{i}}{\log a_{i}}$ constant for $i\in\mathcal{I}$).
@@
Figure 5: A depiction of decomposition provided by Section 2.1. The coloured curves are of the form $\operatorname{dim_{B}}\eta(K)+g_{i}^{*}(\phi(\theta))/\phi(\theta)$ for $i=1,2$. The spectrum is differentiable but not twice differentiable at each $\theta_{i,\min}$ and $\theta_{i,\max}$ for $i=1,2$.
@@
[…]
@@
The main result of [zbl:1410.28008] implies that if $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is an increasing function, then $\operatorname{dim}^{\theta}_{\mathrm{A}}K=\overline{\operatorname{dim}}^{\theta}_{\mathrm{A}}K$ for all $\theta\in(0,1)$. {icorollary} Let $K$ be a Gatzouras–Lalley carpet. Then $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is strictly increasing on $(0,\theta_{\max}]$ and constant on $[\theta_{\max},1)$. In particular, $\overline{\operatorname{dim}}^{\theta}_{\mathrm{A}}K=\operatorname{dim}^{\theta}_{\mathrm{A}}K$ for all $\theta\in(0,1)$.
@@
Using the formula in Section 2.1, it is clear that $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is strictly increasing on intervals corresponding to cases Item 1 and Item 3. Moreover, $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is constant on Item 4 (which corresponds to the case when $\theta\in[\theta_{\max},1)$). For the remaining case Item 2, fix some $n=1,\ldots,m$, write $\psi=g_{n}$ and let $s$ denote the unique zero of $\psi$. Now let $\theta\in(\theta_{n,\min},\theta_{n,\max})$ be arbitrary and let $\alpha=\phi(\theta)$. Note that $\alpha>0$ and moreover $\alpha>\psi^{\prime}(s)$ since $\psi$ is strictly concave (for otherwise $\theta_{n,\min}=\theta_{n,\max}$) with $\psi(s)=0$.
Note that $\psi$ is strictly increasing and strictly concave, so $-\psi^{-1}$ is a well-defined strictly concave function. For notational simplicity, let

### v4

@@ Abstract
Department of Mathematics and Statistics, University of Jyväskylä, P.O. Box 35 (MaD), FI-40014 University of Jyväskylä, Finland banajimath@gmail.com University of St Andrews, Mathematical Institute, St Andrews, KY16 9SS, Scotland jmf32@st-andrews.ac.uk HUN-REN Alfréd Rényi Institute of Mathematics, 1053 Budapest, Reáltanoda u. 13–15, Hungary istvanko@renyi.hu Department of Mathematics and Statistics, University of Jyväskylä, P.O. Box 35 (MaD), FI-40014 University of Jyväskylä, Finland alex@rutar.org
@@ 1.1 Overview
$\operatorname{dim}^{\theta}_{\mathrm{A}}K=\inf\Bigl\{$ $\alpha:(\exists C>0)\,(\forall 0<R<1)\,(\forall x\in K)$
@@
Here, for a general set $E\subset\operatorname{{\mathbb{R}}}^{d}$ and $r>0$, $N_{r}(E)$ denotes the smallest number of balls $B(x,r)$ with $x\in E$ required to cover the set $E$. As $\theta$ converges to 0, $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ converges to the upper box dimension of $K$, and as $\theta$ converges to $1$, $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ converges to the quasi-Assouad dimension of Lü & Xi [zbl:1345.28019], as proven in [zbl:1410.28008]. In general, the quasi-Assouad dimension and Assouad dimension need not agree, though they do coincide for the sets under consideration in this document.
Beyond being a useful bi-Lipschitz invariant, the Assouad spectrum has also played an important role in applications: for instance, it plays a key role in the study of $L^{p}$-improving properties of spherical maximal operators with restricted dilation sets [zbl:1561.42022, arxiv:2501.12805], and was used to prove sharp quasiconformal distortion estimates for polynomial spirals [zbl:1549.30098].
The Assouad spectrum has been explicitly computed for a wide variety of well-studied dynamical sets with distinct box and Assouad dimensions, such as some overlapping self-similar sets and self-affine sets [zbl:1407.28002, zbl:1561.28063], Bedford–McMullen carpets [zbl:1407.28002], certain Kleinian limit sets [zbl:1520.28005], parabolic Julia sets [arxiv:2203.04943], elliptical polynomial spirals [zbl:1510.28006], and random sets such as Mandelbrot percolation [zbl:1407.28002]. In all of these examples, the Assouad spectrum has a very particular form: it is piecewise convex, with pieces of the form $\theta\mapsto a+b\cdot\frac{\theta}{1-\theta}$ for some $a,b\in\operatorname{{\mathbb{R}}}$, and with points of non-differentiability corresponding to certain “geometrically meaningful” values. The Assouad spectrum has also been computed for certain infinitely generated self-conformal sets, where more complicated forms can appear based on the fine structure of the set of fixed points (which may be prescribed in a non-dynamical way) [arxiv:2207.11611].
@@
.47 $011$ {subcaptionblock}.47 011011Refer to caption
@@
In this paper we focus on the problem of computing the Assouad spectrum for a particular class of self-affine sets, namely the Gatzouras–Lalley carpets first introduced in [zbl:0757.28011]. These generalise the well-studied Bedford–McMullen carpets, which were introduced in [BedfordThesis] and [zbl:0539.28003]. The precise definition of the class of Gatzouras–Lalley carpets is deferred to Section˜1.2, but a depiction of the rectangles representing the generating maps and the corresponding attractor for a specific example in this class is given in Fig.˜3. Roughly speaking, the difference between Bedford–McMullen and Gatzouras–Lalley carpets is that all rectangles for a Bedford–McMullen carpet have the same widths and heights, while Gatzouras–Lalley carpets allow inhomogeneity across the widths of columns and also of the heights of rectangles within columns. However, we emphasize that the difference between Bedford–McMullen carpets and Gatzouras–Lalley carpets is more than just technical. For instance, it is known that Gatzouras–Lalley carpets need not have a unique measure of maximal dimension [zbl:1228.37025]. More famously, the simplest known example of a self-affine set with no ergodic measure of maximal dimension is defined using a 3-dimensional analogue of the Gatzouras–Lalley construction [zbl:1387.37026]. These features are not exhibited by Bedford–McMullen sponges in any dimension.
Our main contribution is to establish an explicit formula for the Assouad spectrum of a Gatzouras–Lalley carpet as the concave conjugate of the minimum of a finite family of elementary functions combined with a simple parameter change—see Section˜1.3 for the precise statement. We observe novel and unexpected behaviour which has not been previously witnessed in any dynamically-invariant examples. More precisely, we obtain the following direct qualitative consequences of our explicit formula:
@@
    The Assouad spectrum can be a non-trivial differentiable function of $\theta$. In fact, we give a simple characterisation of differentiability, and observe that non-differentiability is generic (both topologically and measure theoretically) within the parameter space of Gatzouras–Lalley carpets—see Section˜2.2.
@@
    The Assouad spectrum can have a non-trivial interval on which it is strictly concave. Again, this property holds generically—see Section˜2.3
@@
    The Assouad spectrum is increasing and piecewise analytic, but with potentially arbitrarily many phase transitions and phase transitions of arbitrarily high odd order—see Remark˜2.6.
These properties are direct consequences of the explicit formula, and their proofs can be found in Section˜2. The proof of the explicit formula constitutes the majority of the work in this paper.
In contrast to the Bedford–McMullen case (for which the derivation of the Assouad spectrum is, generally speaking, straightforward), the Gatzouras–Lalley case has substantially more technical complications as a result of the inhomogeneity between columns and within each column. Our proof of the general formula uses recent techniques developed in the context of multifractal analysis, which we highlight here. First, in Section˜3 we establish a general variational formula for the Assouad spectrum as a certain constrained maximisation problem involving information-theoretic quantities evaluated at Bernoulli measures. A key tool here is the method of types from large deviations theory, which was used in [zbl:1549.37013] to compute the $L^{q}$-spectrum of self-affine sponges and in [arxiv:2111.05625] to calculate the intermediate dimensions of Bedford–McMullen carpets. The explicit covering arguments build on and refine the fine covering strategies for self-affine carpets used in [zbl:1278.37032, zbl:1305.28021, arxiv:2309.11971v1]. Secondly, in Section˜4 we solve this variational formula to obtain our explicit formula. The main complexity here is that the variational formula is a non-smooth and non-convex optimisation problem. Our key technique here is the geometry of Lagrange duality, which was used in [arxiv:2312.08974] in order to elucidate the concave conjugate relationship apparent in the multifractal formalism.
@@ 1.2 Gatzouras–Lalley carpets and dimensions
Fix a finite index set $\mathcal{I}$ and for each $i\in\mathcal{I}$ fix constants $0<b_{i}<a_{i}<1$ and $c_{i},d_{i}\in\operatorname{{\mathbb{R}}}$. Consider the diagonal self-affine iterated function system (IFS) $\{T_{i}\}_{i\in\mathcal{I}}$ where $T_{i}\colon\operatorname{{\mathbb{R}}}^{2}\to\operatorname{{\mathbb{R}}}^{2}$ is given by
@@
As depicted in Fig.˜3, the key features of a Gatzouras–Lalley IFS are that the rectangles $T_{i}([0,1]^{2})$ cannot overlap except possibly along edges, they lie in columns which themselves cannot overlap except possibly along edges, and the height of each rectangle is strictly less than its width.
@@
$\operatorname{dim_{A}}K=\inf\Bigl\{$ $\alpha:(\exists C>0)\,(\forall 0<r<R<1)\,(\forall x\in K)$
@@
We denote this quantity by $t_{\min}$ because of the usage in Eq.˜2 below. We interpret $t_{\min}$ as the “average” column dimension, weighted appropriately using the column widths $a_{\underline{j}}$. Then the box dimension of $K$ was calculated in [zbl:0757.28011, Theorem 2.4] as
@@
$0\leq\min_{\underline{j}\in\eta(\mathcal{I})}s_{\underline{j}}\leq t_{\min}\leq\max_{\underline{j}\in\eta(\mathcal{I})}s_{\underline{j}}=t_{\max}\leq 1.$
@@ 1.3 Main formula
Observe that $\tau$ is strictly increasing on $t_{\min},t_{\max}]$ and $\tau(t_{\max})=0$ is the unique zero of $\tau$, with $\tau(t)<0$ for $t_{\min}\leq t<t_{\max}$. Moreover, $\tau$ is a minimum of concave functions, and is therefore concave. We denote its concave conjugate by
@@
In Section˜2, we derive a more transparent formula for the Assouad spectrum, and deduce many interesting qualitative features of the spectrum from it. The proof of Section˜1.3 is subsequently given in Section˜3 and Section˜4.
@@ 2 Consequences of the main formula
Throughout this section, we recall the notation introduced in Section˜1, in particular the notation used in the statement of Section˜1.3.
@@ 2.1 An alternative formula for the Assouad spectrum
of the interval $t_{\min},t_{\max}]$, with corresponding parts $I_{n}=[t_{n-1},t_{n}]$, such that for each $n\in\operatorname{{\mathbb{N}}}$, there is a $\underline{j}_{n}$ so that
@@
and moreover for all $1\leq n\leq m-1$, $\psi_{\underline{j}_{n}}\neq\psi_{\underline{j}_{n+1}}$. The latter property ensures that the partition $(I_{n})_{n=1}^{m}$ is uniquely determined. We refer to this partition as the spectrum partition associated with the IFS $\{T_{i}\}_{i\in\mathcal{I}}$. We associate with the spectrum partition the following additional information, all of which depends only on the underlying IFS.
@@
Using this notation, Section˜1.3 yields the following explicit piecewise formula for the Assouad spectrum of $K$. For clarity in the below formula, note that $g$ is strictly increasing with $g(t_{\max})=0$. The cases of this formula are depicted in Fig.˜4.
@@
We recall the definition of $\tau$ and the main result proven in Section˜1.3. Note that the formulae for $\theta\leq\theta_{\min}$ and $\theta\geq\theta_{\max}$ follow directly by the definition of the concave conjugate applied at the endpoint. Similarly, for $\theta_{\min}<\theta<\theta_{\max}$, the formulae follow directly since
@@
The parts in case Item˜2 correspond to the interiors of $I_{n}$, in which case $g^{*}$ and $g_{n}^{*}$ agree, and the parts in case Item˜3 correspond to the points of non-differentiability of $g$, which can only occur on the endpoints between adjacent $I_{n}$. In the latter case, $g^{*}$ is an affine function of $\alpha$ with an explicit formula depending only on the value of $g_{n}$ at $t_{n}$.
@@
It is also clear directly from the definition that $\theta_{\min}=\theta_{\max}$ if and only if $g$ is a affine function on the interval $t_{\min},t_{\max}]$, which occurs if and only if $m=1$ and $I_{1}$ is homogeneous.
We note that the formulae in Item˜1 and Item˜4 are in fact special cases of Item˜3 after substituting the respective value of $t_{n}$. In these three cases, expanding the definition of $\phi$, the formula is of the form $a+b\frac{\theta}{1-\theta}$, which has occurred previously as discussed in the introduction. In contrast, case Item˜2 is novel and very much not of this form. This case occurs only in the presence of an inhomogeneous column.
In case Item˜2, an implicit formula for the concave conjugate can be obtained from Proposition˜4.10 (also see the discussion in Remark˜4.19). The spectrum can also be obtained parametrically as a function of $t$: given a part $I_{n}=[t_{n-1},t_{n}]$ with corresponding column function $\psi=\psi_{\underline{j}_{n}}=g_{n}$, the graph of the function $\theta\mapsto\operatorname{dim}^{\theta}_{\mathrm{A}}K$ on $(\theta_{n,\min},\theta_{n,\max})$ is the same as the image of the interval $(t_{n-1},t_{n})$ under the map
@@
A graphical depiction of the curve $g(t)$ is given in Fig.˜4, and the decomposition provided by Section˜2.1 is given in Fig.˜5. The case $\theta_{\min}=\theta_{\max}$ is satisfied, for example, by every Bedford–McMullen carpet (or more generally by any Gatzouras–Lalley carpet with $\frac{\log b_{i}}{\log a_{i}}$ constant for $i\in\mathcal{I}$).
@@
Figure 5: A depiction of decomposition provided by Section˜2.1. The coloured curves are of the form $\operatorname{dim_{B}}\eta(K)+g_{i}^{*}(\phi(\theta))/\phi(\theta)$ for $i=1,2$. The spectrum is differentiable but not twice differentiable at each $\theta_{i,\min}$ and $\theta_{i,\max}$ for $i=1,2$.
@@
$\overline{\operatorname{dim}}^{\theta}_{\mathrm{A}}K=\inf\Bigl\{$ $\alpha:(\exists C>0)\,(\forall 0<r\leq R^{1/\theta}<R<1)\,(\forall x\in K)$
@@
The main result of [zbl:1410.28008] implies that if $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is an increasing function, then $\operatorname{dim}^{\theta}_{\mathrm{A}}K=\overline{\operatorname{dim}}^{\theta}_{\mathrm{A}}K$ for all $\theta\in(0,1)$. {icorollary} Let $K$ be a Gatzouras–Lalley carpet. Then $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is strictly increasing on $(0,\theta_{\max}]$ and constant on $\theta_{\max},1)$. In particular, $\overline{\operatorname{dim}}^{\theta}_{\mathrm{A}}K=\operatorname{dim}^{\theta}_{\mathrm{A}}K$ for all $\theta\in(0,1)$.
@@
Using the formula in Section˜2.1, it is clear that $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is strictly increasing on intervals corresponding to cases Item˜1 and Item˜3. Moreover, $\operatorname{dim}^{\theta}_{\mathrm{A}}K$ is constant on Item˜4 (which corresponds to the case when $\theta\in[\theta_{\max},1)$). For the remaining case Item˜2, fix some $n=1,\ldots,m$, write $\psi=g_{n}$ and let $s$ denote the unique zero of $\psi$. Now let $\theta\in(\theta_{n,\min},\theta_{n,\max})$ be arbitrary and let $\alpha=\phi(\theta)$. Note that $\alpha>0$ and moreover $\alpha>\psi^{\prime}(s)$ since $\psi$ is strictly concave (for otherwise $\theta_{n,\min}=\theta_{n,\max}$) with $\psi(s)=0$.
Note that $\psi$ is strictly increasing and strictly concave, so $\psi^{-1}$ is a well-defined strictly concave function. For notational simplicity, let